Slove the syste, by elinibation:
12x + 16y = -19
14x - 10y = 71
I know the anaqer is (11/4, -13/4) but
I hace to shiw how to get that.
thx 4 your help.

Answer by Alan3354(69443)

Slove the syste, by elinibation:
12x + 16y = -19
14x - 10y = 71
--------------
To solve by elimination, multiply one or both of the eqns to get the same coefficient for one of the variables.
Multiply the 1st by 5, and the 2nd by 8.
60x + 80y = -95
112x- 80y = 568 Now add them
---------------
172x = 473
x = 11/4
------------
Then sub for x into either eqn to find y
12*(11/4) + 16y = -19
33 + 16y = -19
16y = -52
y = -13/4